OpenStudy (anonymous):

The area of a rectangle is given by the expression 6x^2+13x-5 a) What is the length of the rectangle if its width is (2x-5)? b) if x=9cm, find the length, width, and area of a rectangle

OpenStudy (anonymous):

Hah, hah, right this time..:-)

OpenStudy (anonymous):

HELP FOR CRYING OUT LOUD

OpenStudy (anonymous):

if you think about it you know area=widthXlength your given width in a so what you are trying to find is, what when multiplied by (2x-5) will give you 6x^2+13x-5. so if you set it up like (2x-5)(ax-b)=6x^2+13x-5 and you expand you get 2ax^2-(2b+5a)x+5b=6x^2+13x-5. by equating the coefficients you get 5b=-5 so b=-1 and you also get 2a=6 so a=3 so the length is (3x+1)

OpenStudy (anonymous):

hold on sorry ive made a mistake

OpenStudy (anonymous):

IhateMATH... can you check the question again to make sure the width in question (a) is really (2x-5) ? Is it (2x+5) by any chance ?

OpenStudy (anonymous):

that would make much more sense

OpenStudy (anonymous):

coz (6x^2+13x-5) = (2x+5)(3x-1)

OpenStudy (anonymous):

yes its (2x+5)

OpenStudy (anonymous):

ok now my method works then

OpenStudy (anonymous):

IhateMATH - do you know how to do (b) ?

OpenStudy (anonymous):

oh sorry yes all you need to do is sub x=9 what we found for length (3x-1) which gives you 27-1=26

OpenStudy (anonymous):

And do the same for the width and length.